A mother is 25 years older than her daughter. In 10 years, she will be twice as old as her daughter. Determine their present age
Found 2 solutions by Cromlix, Alan3354:
Answer by Cromlix(4381)

Hi there,
A mother is 25 years older than her daughter
M = 25 + D
In 10 years, she will be twice as old as her daughter.
M + 10 = 2(D + 10)
Replace M with 25 + D
25 + D + 10 = 2(D + 10)
25 + D + 10 = 2D + 20
Collect like terms
D - 2D = 20 - 25 - 10
-D = -15
Multiply both sides by -1
D = 15
Therefore present ages:
Daughter = 15 yrs old
Mother = 40 years old.
Hope this helps:-)

Answer by Alan3354(69443)

A mother is 25 years older than her daughter. In 10 years, she will be twice as old as her daughter. Determine their present age
--------
The difference = 25 years
When she's 2x as old --> 25 & 50
Present ages = 15 & 40